# Some lemmas on spectral radius of graphs: including an application

### v2

@@ 2 Preliminaries
The following fact can be proved by induction on $\ell$.
Fact 1. For integers $n\geq 4\ell\geq 0$, we have
$\binom{n}{0}+\binom{n}{1}+\cdots+\binom{n}{\ell}\leq\frac{5}{4}\frac{n^{\ell}}{\ell!}$
@@
Let $G$ be a path of order $n\geq 3$. Let $u,v$ be two vertices of $G$ such that $d_{G}(u,v)=\ell$, where $2\leq\ell\leq n-1$. Then $\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{10x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4})$ for any $x\geq\max\left\{\sqrt{n},5\right\}$.
@@
Claim 1. $w^{a}_{uv}(z_{i})\leq w^{a}_{uv}(u_{i})+w^{a}_{uv}(v_{\ell-i})$ for any $1\leq i\leq\ell-1$ and $a\geq 1$.
Proof of Claim 1. Let $a\geq 1$ be given. Clearly, $w^{a}_{uv}(z_{i})=0$ for $1\leq a\leq\ell-1$. Now assume $a\geq\ell$. Let $1\leq i_{0}\leq\ell-1$ be fixed. We will show that
$w^{a}_{uv}(z_{i_{0}})\leq w^{a}_{uv}(u_{i_{0}})+w^{a}_{uv}(v_{\ell-i_{0}}).$
Let $X$ be the set of walks $Q$ starting at $z_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $u$ first before $v$. Let $Y$ be the set of walks $Q$ starting at $z_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $v$ first before $u$. Clearly, $w^{a}_{uv}(z_{i_{0}})=|X|+|Y|$.
Now we show that $|X|\leq w^{a}_{uv}(u_{i_{0}})$ and $|Y|\leq w^{a}_{uv}(v_{\ell-i_{0}})$. We show the first one. Let $D$ be the set of walks $Q$ starting at $u_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $u,v$. We can define an injective map $\phi$ between $X$ and $D$. For any $Q\in X$, since $Q$ crosses $u$ first, we can denote
@@
It is easy to see that $\phi$ is injective. Hence, $|X|\leq|D|=w^{a}_{uv}(u_{i_{0}})$. Similarly, we can show that $|Y|\leq w^{a}_{uv}(v_{\ell-i_{0}})$. Then $w^{a}_{uv}(z_{i_{0}})=|X|+|Y|\leq w^{a}_{uv}(u_{i_{0}})+w^{a}_{uv}(v_{\ell-i_{0}})$. This finishes the proof of Claim 1. $\Box$
@@
$\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(z_{i})\leq\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(v_{i})$
for any $a\geq 1$. Clearly, $w^{a}_{uv}(u_{i})=w^{a}_{uv}(v_{i})=0$ for any $i\geq\ell+1$ and $a\leq 2\ell$. Thus, for any $a\leq 2\ell$,
$W^{a}_{u,v}(P_{\infty})$ $=w^{a}_{uv}(u)+w^{a}_{uv}(v)+\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(z_{i})+\sum_{1\leq i\leq\ell}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\ell}w^{a}_{uv}(v_{i})$
$\leq 2(w^{a}_{uv}(u)+w^{a}_{uv}(v)+\sum_{1\leq i\leq\ell}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\ell}w^{a}_{uv}(v_{i})).$
By symmetry, $w^{a}_{uv}(u)=w^{a}_{uv}(v)$ and $w^{a}_{uv}(u_{i})=w^{a}_{uv}(v_{i})$ for any $a\geq 1$ and $i\geq 1$. Thus, for any $a\leq 2\ell$,
$W^{a}_{u,v}(P_{\infty})\leq 4(w^{a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{a}_{uv}(u_{i})).$
Since $P_{\infty}$ is a path, each walk starting at $u_{i}$ corresponds to a $\pm$”-series:
@@
(For example, $u_{1}+-++$-- corresponds to the walk $u_{1}uu_{1}uz_{1}uu_{1}$.) For any $0\leq a\leq\ell$ and $1\leq i\leq a$, in the corresponding $\pm$”-series to the walks of length $\ell+a$ starting at $u_{i}$ and crossing $u$ and $v$, there are at most $\lfloor\frac{a-i}{2}\rfloor$ symbols ”-” (in the all $\ell+a$ symbols $\pm$”). Thus, using Fact 1 as $\ell+a\geq 4\lfloor\frac{a-i}{2}\rfloor$,
$w^{\ell+a}_{uv}(u_{i})\leq\binom{\ell+a}{0}+\binom{\ell+a}{1}+\cdots+\binom{\ell+a}{\lfloor\frac{a-i}{2}\rfloor}\leq\frac{5}{4}\frac{(\ell+a)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}.$
Hence, $w^{\ell+a}_{uv}(u_{i})\leq\frac{5}{4}\frac{(2\ell)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}$ for any $0\leq a\leq\ell$. It follows that, for any $0\leq a\leq\ell$,
$w^{\ell+a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{\ell+a}_{uv}(u_{i})\leq 2\times\frac{5}{4}\times(1+2\ell+\frac{(2\ell)^{2}}{2!}+\cdots+\frac{(2\ell)^{\lfloor\frac{a}{2}\rfloor}}{\lfloor\frac{a}{2}\rfloor!})=\frac{5}{2}\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}.$
@@
$\sum_{0\leq a\leq\ell}\frac{W^{\ell+a}_{u,v}(P_{\infty})}{x^{\ell+a}}$ $\leq\sum_{0\leq a\leq\ell}\frac{4}{x^{\ell+a}}(w^{\ell+a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{\ell+a}_{uv}(u_{i}))$
$\leq\frac{10}{x^{\ell}}\sum_{0\leq a\leq\ell}\frac{1}{x^{a}}\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}$
$=\frac{10}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\sum_{2i\leq a\leq\ell}\frac{1}{x^{a}}$
$\leq\frac{10}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\frac{1}{x^{2i}}\frac{x}{x-1}$
$\leq\frac{10}{x^{\ell}}\frac{x}{x-1}\sum^{\infty}_{i=0}\frac{1}{i!}(\frac{2\ell}{x^{2}})^{i}$
$=\frac{10}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
@@
$\sum_{1\leq s\leq 2\ell}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\sum_{0\leq a\leq\ell}\frac{W^{\ell+a}_{u,v}(P_{\infty})}{x^{\ell+a}}\leq\frac{10}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
Since $G$ has maximum degree 2, we see $W^{a}_{u,v}(G)\leq 2^{a}n$ for any $a\geq 1$. Noting $s-3\leq 2(s-\ell-2)$ for $s\geq 2\ell+1$, we have
$\sum_{s\geq 2\ell+1}\frac{W^{s}_{u,v}(G)}{x^{s}}$ $\leq\sum_{s\geq 2\ell+1}\frac{2^{s}n}{x^{s}}$
$\leq\frac{8}{x^{\ell}}\sum_{s\geq 2\ell+1}\frac{2^{s-3}}{x^{s-\ell-2}}~~~(using~x\geq\sqrt{n})$
$\leq\frac{8}{x^{\ell}}\sum_{s\geq 2\ell+1}(\frac{4}{x})^{s-\ell-2}~~~(using~s-3\leq 2(s-\ell-2))$
$\leq\frac{8}{x^{\ell}}\sum_{j\geq 1}(\frac{4}{x})^{j}~~~(using~\ell\geq 2)$
$=\frac{1}{x^{\ell}}\frac{32}{x-4}~~~(using~x\geq 5).$
@@
$\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{10x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4}).$
@@
$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 4}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 4}n_{i}}\geq\max\left\{\sqrt{130},\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}.$
@@
$\sum^{\infty}_{i=1}\frac{W^{i}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{1}{\rho_{1}^{n_{4}+1}}(\frac{10\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
Since $\rho_{1}\geq\max\left\{\sqrt{130},\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}\geq\max\left\{10,\sqrt{4n_{4}+2}\right\}$, we have $e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}=e^{\frac{(2n_{4}+1)+1}{\rho_{1}^{2}}}\leq e^{0.51}<1.67$. Thus,
@@
$\leq\frac{2}{\rho_{1}^{n_{4}+1}}(\frac{5\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{16}{\rho_{1}-4}-\rho_{1}-1)$

### v3

@@ 1 Introduction
Note. In this third version, we correct some calculation errors in Lemma 3.3.
@@ 2 Preliminaries
The following result will be used.
@@
For integers $n\geq 6m\geq 0$, we have
$\binom{n}{0}4^{m}+\binom{n}{1}4^{m-1}+\cdots+\binom{n}{m}4^{0}\leq\frac{5}{2}\frac{n^{m}}{m!}$
Proof: To prove the lemma, it is equivalent to show
$\binom{n}{0}+\binom{n}{1}\frac{1}{4}+\binom{n}{2}\frac{1}{4^{2}}+\cdots+\binom{n}{m}\frac{1}{4^{m}}\leq\frac{5}{2}\frac{n^{m}}{4^{m}m!}.$
We will prove this inequality by induction on $m$. It is easy to check for $m=0,1$. When $m=2$, we need to show $\binom{n}{0}+\binom{n}{1}\frac{1}{4^{1}}+\binom{n}{2}\frac{1}{4^{2}}\leq\frac{5}{2}\frac{n^{2}}{4^{2}2}$, where $n\geq 12$. This is equivalent to show $\frac{3}{2}n^{2}-3n-16\geq 0$, which is clearly true as $n\geq 12$. When $m=3$, we need to show $\binom{n}{0}+\binom{n}{1}\frac{1}{4^{1}}+\binom{n}{2}\frac{1}{4^{2}}+\binom{n}{3}\frac{1}{4^{3}}\leq\frac{5}{2}\frac{n^{3}}{4^{3}6}$, where $n\geq 18$. This is equivalent to show $3n^{3}-18n^{2}-172n-768\geq 0$, which is clearly true as $n\geq 18$.
Now we suppose the inequality holds for $m=k$, where $k\geq 3$. We will prove the inequality for $m=k+1$. Now let $m=k+1$, implying $n\geq 6(k+1)$. Since the inequality holds for $m=k$, we have
$\binom{n}{0}+\binom{n}{1}\frac{1}{4}+\binom{n}{2}\frac{1}{4^{2}}+\cdots+\binom{n}{k}\frac{1}{4^{k}}\leq\frac{5}{2}\frac{n^{k}}{4^{k}k!}.$
Then
$\binom{n}{0}+\binom{n}{1}\frac{1}{4}+\binom{n}{2}\frac{1}{4^{2}}+\cdots+\binom{n}{k}\frac{1}{4^{k}}+\binom{n}{k+1}\frac{1}{4^{k+1}}\leq\binom{n}{k+1}\frac{1}{4^{k+1}}+\frac{5}{2}\frac{n^{k}}{4^{k}k!}.$
To prove the inequality for $m=k+1$, it suffices to show
$\binom{n}{k+1}\frac{1}{4^{k+1}}+\frac{5}{2}\frac{n^{k}}{4^{k}k!}\leq\frac{5}{2}\frac{n^{k+1}}{4^{k+1}(k+1)!}.$
This is equivalent to show
$(n-1)(n-2)\cdots(n-k)\leq\frac{5}{2}n^{k-1}(n-4k-4).$
It suffices to show
$(n-k+1)(n-k)\leq\frac{5}{2}n(n-4k-4).$
This is equivalent to show
$n(3n-16k-22)\geq 2k^{2}-2k,$
which is clearly true as $n\geq 6k+6$ and $k\geq 3$. Thus, the inequality also holds for $m=k+1$. By induction on $m$, this completes the proof. $\Box$
Lemma 3.3
Let $G$ be a path of order $n\geq 3$. Let $u,v$ be two vertices of $G$ such that $d_{G}(u,v)=\ell$, where $2\leq\ell\leq n-1$. Then
$\sum_{a\geq 1}\frac{W^{a}_{u,v}(G)}{x^{a}}\leq\frac{1}{x^{\ell}}(\frac{30x}{x-1}e^{\frac{\lfloor\frac{3\ell}{2}\rfloor}{x^{2}}}+\frac{32x}{x-8})$
for any $x\geq\max\left\{\sqrt{n},9\right\}$.
@@
Claim 1. Let $1\leq a\leq\frac{3\ell}{2}$ be an integer. Then $w^{a}_{uv}(z_{i})\leq w^{a}_{uv}(u_{i})$ and $w^{a}_{uv}(z_{\ell-i})\leq w^{a}_{uv}(v_{i})$ for any $1\leq i<\frac{\ell}{2}$. Moreover, $w^{a}_{uv}(z_{\frac{\ell}{2}})\leq w^{a}_{uv}(u_{\frac{\ell}{2}})+w^{a}_{uv}(v_{\frac{\ell}{2}})$ for even $\ell$.
Proof of Claim 1. Clearly, $w^{a}_{uv}(z_{i})=0$ for $1\leq a\leq\ell-1$. Now assume $\ell\leq a\leq\frac{3\ell}{2}$. Let $1\leq i_{0}<\frac{\ell}{2}$ be fixed. We will show that
$w^{a}_{uv}(z_{i_{0}})\leq w^{a}_{uv}(u_{i_{0}}).$
Let $X$ be the set of walks $Q$ starting at $z_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $u$ and $v$. Clearly, $w^{a}_{uv}(z_{i_{0}})=|X|$. Note that $1\leq i_{0}<\frac{\ell}{2}$ and $a\leq\frac{3\ell}{2}$. Each $Q\in X$ crosses $u$ first.
Now we show that $|X|\leq w^{a}_{uv}(u_{i_{0}})$. Let $D$ be the set of walks $Q$ starting at $u_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $u$ and $v$. Clearly, $w^{a}_{uv}(u_{i_{0}})=|D|$. We can define an injective map $\phi$ between $X$ and $D$. For any $Q\in X$, since $Q$ crosses $u$ first, we can denote
@@
It is easy to see that $\phi$ is injective. Hence, $|X|\leq|D|=w^{a}_{uv}(u_{i_{0}})$. That is, $w^{a}_{uv}(z_{i_{0}})\leq w^{a}_{uv}(u_{i_{0}})$. Similarly, we can show that $w^{a}_{uv}(z_{\ell-i_{0}})\leq w^{a}_{uv}(v_{i_{0}})$.
Now assume that $\ell\geq 2$ is even. Let $Y_{u}$ be the set of walks $Q$ starting at $z_{\frac{\ell}{2}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $u$ first before $v$. Let $Y_{v}$ be the set of walks $Q$ starting at $z_{\frac{\ell}{2}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $v$ first before $u$. Clearly, $w^{a}_{uv}(z_{\frac{\ell}{2}})=|Y_{u}|+|Y_{v}|$. Similar to the above discussion, we can show that $|Y_{u}|\leq w^{a}_{uv}(u_{\frac{\ell}{2}})$ and $|Y_{v}|\leq w^{a}_{uv}(v_{\frac{\ell}{2}})$. Hence $w^{a}_{uv}(z_{\frac{\ell}{2}})\leq w^{a}_{uv}(u_{\frac{\ell}{2}})+w^{a}_{uv}(v_{\frac{\ell}{2}})$. This finishes the proof of Claim 1. $\Box$
@@
$\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(z_{i})\leq\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(v_{i})$
for any $1\leq a\leq\frac{3}{2}\ell$. Clearly, $w^{a}_{uv}(u_{i})=w^{a}_{uv}(v_{i})=0$ for any $i>\frac{\ell}{2}$ and $a\leq\frac{3}{2}\ell$. Thus, for any $a\leq\frac{3}{2}\ell$,
$W^{a}_{u,v}(P_{\infty})$ $=w^{a}_{uv}(u)+w^{a}_{uv}(v)+\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(z_{i})+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(v_{i})$
$\leq 2\left(w^{a}_{uv}(u)+w^{a}_{uv}(v)+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(v_{i})\right).$
By symmetry, $w^{a}_{uv}(u)=w^{a}_{uv}(v)$ and $w^{a}_{uv}(u_{i})=w^{a}_{uv}(v_{i})$ for any $a\geq 1$ and $i\geq 1$. Thus, for any $a\leq\frac{3\ell}{2}$,
$W^{a}_{u,v}(P_{\infty})\leq 4(w^{a}_{uv}(u)+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(u_{i})).$
Claim 2. Let $a,i$ be given integers with $\ell\leq a\leq\frac{3\ell}{2}$ and $1\leq i\leq\frac{\ell}{2}$. Then
$w^{a}_{uv}(u_{i})\leq\frac{5}{2}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{\lfloor\frac{a-\ell-i}{2}\rfloor}}{\lfloor\frac{a-\ell-i}{2}\rfloor!}$
for even $a-\ell-i$, and
$w^{a}_{uv}(u_{i})\leq\frac{5(\lfloor\frac{3\ell}{2}\rfloor)^{\lfloor\frac{a-\ell-i}{2}\rfloor}}{\lfloor\frac{a-\ell-i}{2}\rfloor!}$
for odd $a-\ell-i$.
Proof of Claim 2. Since $P_{\infty}$ is a path, each walk starting at $u_{i}$ corresponds to a $\pm$”-series:
@@
(For example, $u_{1}+-++$-- corresponds to the walk $u_{1}uu_{1}uz_{1}uu_{1}$.) Recall that $\ell\leq a\leq\frac{3\ell}{2}$ and $1\leq i\leq\frac{\ell}{2}$. Note that the distance between $u_{i}$ and $v$ is $\ell+i$. For any $\ell+i\leq j\leq a$, let $X_{j}$ be the set of walks $Q$ starting at $u_{i}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $v$ at step $j$ (with smallest value). That is, $Q=u_{i}g_{1}g_{2}\cdots g_{a}$, where $g_{j}=v$ and $g_{k}\neq v$ for any $1\leq k\leq j-1$. Clearly, $j\equiv\ell+i(\hbox{\rm mod }~2)$. Moreover,
$w^{a}_{uv}(u_{i})=|X_{\ell+i}|+|X_{\ell+i+2}|+\cdots+|X_{a}|$
for even $a-\ell-i$, and
$w^{a}_{uv}(u_{i})=|X_{\ell+i}|+|X_{\ell+i+2}|+\cdots+|X_{a-1}|$
for odd $a-\ell-i$.
For any $\ell+i\leq j\leq a$ and $j\equiv\ell+i(\hbox{\rm mod }~2)$, in the corresponding $\pm$”-series to walks $Q\in X_{j}$, there are exactly $\lfloor\frac{j-\ell-i}{2}\rfloor$ symbols ”-” in the first $j$ symbols $\pm$”. Hence,
$|X_{j}|\leq\binom{j}{\lfloor\frac{j-\ell-i}{2}\rfloor}2^{a-j}.$
Since $j\leq a\leq\lfloor\frac{3\ell}{2}\rfloor$, we have
$|X_{j}|\leq\binom{\lfloor\frac{3\ell}{2}\rfloor}{\lfloor\frac{j-\ell-i}{2}\rfloor}2^{a-j}.$
Let $m=\lfloor\frac{a-\ell-i}{2}\rfloor$. Clearly, $m\leq\frac{a-\ell-i}{2}\leq\frac{\ell}{4}$ as $a\leq\frac{3\ell}{2}$. Then $6m\leq\frac{3\ell}{2}$, implying $6m\leq\lfloor\frac{3\ell}{2}\rfloor$.
For even $a-\ell-i$, using Lemma 3.2 as $6m\leq\lfloor\frac{3\ell}{2}\rfloor$,
$w^{a}_{uv}(u_{i})$ $=|X_{\ell+i}|+|X_{\ell+i+2}|+\cdots+|X_{a}|$
$\leq\binom{\lfloor\frac{3\ell}{2}\rfloor}{0}2^{a-\ell-i}+\binom{\lfloor\frac{3\ell}{2}\rfloor}{1}2^{a-\ell-i-2}+\binom{\lfloor\frac{3\ell}{2}\rfloor}{2}2^{a-\ell-i-4}+\cdots+\binom{\lfloor\frac{3\ell}{2}\rfloor}{\frac{a-\ell-i}{2}}2^{0}$
$=\binom{\lfloor\frac{3\ell}{2}\rfloor}{0}4^{m}+\binom{\lfloor\frac{3\ell}{2}\rfloor}{1}4^{m-1}+\cdots+\binom{\lfloor\frac{3\ell}{2}\rfloor}{m}4^{0}$
$\leq\frac{5}{2}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{m}}{m!}$
$=\frac{5}{2}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{\lfloor\frac{a-\ell-i}{2}\rfloor}}{\lfloor\frac{a-\ell-i}{2}\rfloor!}.$
For odd $a-\ell-i$, we have $m=\frac{a-1-\ell-i}{2}$. Using Lemma 3.2 as $6m\leq\lfloor\frac{3\ell}{2}\rfloor$,
$w^{a}_{uv}(u_{i})$ $=|X_{\ell+i}|+|X_{\ell+i+2}|+\cdots+|X_{a-1}|$
$\leq\binom{\lfloor\frac{3\ell}{2}\rfloor}{0}2^{a-\ell-i}+\binom{\lfloor\frac{3\ell}{2}\rfloor}{1}2^{a-\ell-i-2}+\binom{\lfloor\frac{3\ell}{2}\rfloor}{2}2^{a-\ell-i-4}+\cdots+\binom{\lfloor\frac{3\ell}{2}\rfloor}{\lfloor\frac{a-\ell-i}{2}\rfloor}2^{1}$
$=2\binom{\lfloor\frac{3\ell}{2}\rfloor}{0}2^{a-1-\ell-i}+2\binom{\lfloor\frac{3\ell}{2}\rfloor}{1}2^{a-1-\ell-i-2}+2\binom{\lfloor\frac{3\ell}{2}\rfloor}{2}2^{a-1-\ell-i-4}+\cdots+2\binom{\lfloor\frac{3\ell}{2}\rfloor}{\lfloor\frac{a-\ell-i}{2}\rfloor}2^{0}$
$=2\binom{\lfloor\frac{3\ell}{2}\rfloor}{0}4^{m}+2\binom{\lfloor\frac{3\ell}{2}\rfloor}{1}4^{m-1}+\cdots+2\binom{\lfloor\frac{3\ell}{2}\rfloor}{m}4^{0}$
$\leq\frac{5(\lfloor\frac{3\ell}{2}\rfloor)^{m}}{m!}$
$=\frac{5(\lfloor\frac{3\ell}{2}\rfloor)^{\lfloor\frac{a-\ell-i}{2}\rfloor}}{\lfloor\frac{a-\ell-i}{2}\rfloor!}.$
This finishes the proof of Claim 2. $\Box$
Similar to Claim 2, we can show $w^{a}_{uv}(u)\leq\frac{5}{2}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{\lfloor\frac{a-\ell}{2}\rfloor}}{\lfloor\frac{a-\ell}{2}\rfloor!}$ for even $a$-$\ell$, and $w^{a}_{uv}(u)\leq\frac{5(\lfloor\frac{3\ell}{2}\rfloor)^{\lfloor\frac{a-\ell}{2}\rfloor}}{\lfloor\frac{a-\ell}{2}\rfloor!}$ for odd $a$-$\ell$. Clearly, $w^{a}_{uv}(u_{i})=0$ for $i>a$-$\ell$. By Claim 2, for any $\ell\leq a\leq\frac{3\ell}{2}$,
$w^{a}_{uv}(u)+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(u_{i})$
$=w^{a}_{uv}(u)+\sum_{1\leq i\leq a-\ell}w^{a}_{uv}(u_{i})$
$\leq(5+\frac{5}{2})\times(1+\lfloor\frac{3\ell}{2}\rfloor+\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{2}}{2!}+\cdots+\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{\lfloor\frac{a-\ell}{2}\rfloor}}{\lfloor\frac{a-\ell}{2}\rfloor!})$
$=\frac{15}{2}\sum_{0\leq i\leq\frac{a-\ell}{2}}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{i}}{i!}.$
@@
$\sum_{\ell\leq a\leq\frac{3\ell}{2}}\frac{W^{a}_{u,v}(P_{\infty})}{x^{a}}$ $\leq\sum_{\ell\leq a\leq\frac{3\ell}{2}}\frac{4}{x^{a}}(w^{a}_{uv}(u)+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(u_{i}))$
$\leq\frac{30}{x^{\ell}}\sum_{\ell\leq a\leq\frac{3\ell}{2}}\frac{1}{x^{a-\ell}}\sum_{0\leq i\leq\frac{a-\ell}{2}}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{i}}{i!}$
$=\frac{30}{x^{\ell}}\sum_{0\leq i\leq\frac{\ell}{4}}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{i}}{i!}\sum_{\ell+2i\leq a\leq\frac{3\ell}{2}}\frac{1}{x^{a-\ell}}$
$\leq\frac{30}{x^{\ell}}\sum_{0\leq i\leq\frac{\ell}{4}}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{i}}{i!}\frac{1}{x^{2i}}\frac{x}{x-1}$
$\leq\frac{30}{x^{\ell}}\frac{x}{x-1}\sum^{\infty}_{i=0}\frac{1}{i!}(\frac{\lfloor\frac{3\ell}{2}\rfloor}{x^{2}})^{i}$
$=\frac{30}{x^{\ell}}\frac{x}{x-1}e^{\frac{\lfloor\frac{3\ell}{2}\rfloor}{x^{2}}}.$
@@
$\sum_{1\leq a\leq\frac{3\ell}{2}}\frac{W^{a}_{u,v}(G)}{x^{a}}\leq\sum_{\ell\leq a\leq\frac{3\ell}{2}}\frac{W^{a}_{u,v}(P_{\infty})}{x^{a}}\leq\frac{30}{x^{\ell}}\frac{x}{x-1}e^{\frac{\lfloor\frac{3\ell}{2}\rfloor}{x^{2}}}.$
Since $G$ has maximum degree 2, we see $W^{a}_{u,v}(G)\leq 2^{a}n$ for any $a\geq 1$. Note that $a-5\leq 3(a-\ell-2)$ and $a-\ell-2\geq 0$, if $a>\frac{3\ell}{2}$ and $\ell\geq 2$. Thus,
$\sum_{a>\frac{3\ell}{2}}\frac{W^{a}_{u,v}(G)}{x^{a}}$ $\leq\sum_{a>\frac{3\ell}{2}}\frac{2^{a}n}{x^{a}}$
$\leq\frac{32}{x^{\ell}}\sum_{a>\frac{3\ell}{2}}\frac{2^{a-5}}{x^{a-\ell-2}}~~~(using~x\geq\sqrt{n})$
$\leq\frac{32}{x^{\ell}}\sum_{a>\frac{3\ell}{2}}(\frac{8}{x})^{a-\ell-2}~~~(using~a-5\leq 3(a-\ell-2))$
$\leq\frac{32}{x^{\ell}}\sum_{j\geq 0}(\frac{8}{x})^{j}~~~(using~a-\ell-2\geq 0)$
$=\frac{32}{x^{\ell}}\frac{x}{x-8}~~~(using~x\geq 9).$
@@
$\sum_{a\geq 1}\frac{W^{a}_{u,v}(G)}{x^{a}}\leq\frac{1}{x^{\ell}}(\frac{30x}{x-1}e^{\frac{\lfloor\frac{3\ell}{2}\rfloor}{x^{2}}}+\frac{32x}{x-8}).$
@@
$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 4}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 4}n_{i}}\geq\max\left\{\sqrt{2000},\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}.$
@@
$\sum^{\infty}_{i=1}\frac{W^{i}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{1}{\rho_{1}^{n_{4}+1}}(\frac{30\rho_{1}}{\rho_{1}-1}e^{\frac{\frac{3}{2}(n_{4}+1)}{\rho_{1}^{2}}}+\frac{32\rho_{1}}{\rho_{1}-8}).$
Since $\rho_{1}\geq\max\left\{\sqrt{2000},\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}\geq\max\left\{\sqrt{2000},\sqrt{4n_{4}+2}\right\}$, we have $e^{\frac{\frac{3}{2}(n_{4}+1)}{\rho_{1}^{2}}}\leq e^{\frac{\frac{3}{2}(n_{4}+1)}{4n_{4}+2}}\leq e^{0.5}<1.65$. Thus,
@@
$\leq\frac{2}{\rho_{1}^{n_{4}+1}}(\frac{15\rho_{1}}{\rho_{1}-1}e^{\frac{\frac{3}{2}(n_{4}+1)}{\rho_{1}^{2}}}+\frac{16\rho_{1}}{\rho_{1}-8}-\rho_{1}-1)$